### v2

# Find the Leak, Fix the Split: Cluster-Based Method to Prevent Leakage in Video-Derived Datasets

Abstract
We propose a cluster-based frame selection strategy to mitigate information leakage in video-derived frames datasets. By grouping visually similar frames before splitting into training, validation, and test sets, the method produces more representative, balanced, and reliable dataset partitions.
I Introduction
In deep learning research, database construction and data splitting strategies play a crucial role in determining the reliability and generalization capability of trained models. Image-based models can be trained by randomly splitting images into training, validation, and test sets [2]. However, when datasets are derived from video sources, this approach can lead to information leakage due to high spatial and temporal correlations among consecutive frames, e.g., frames recorded in the same background and containing the same objects but in slightly different positions or rotations [4, 2]. As a result, models can achieve inflated performance estimates, particularly in image-based object detection trained on video-derived frames. To address this issue, we propose a cluster-based frame selection strategy that groups visually similar frames before splitting into the dataset partitions (train, validation, test), ensuring that correlated images remain together. The proposed method is simple, scalable, and can be integrated into existing dataset preparation pipelines without modification of the training process. This approach can improve the fairness of dataset partitioning and improve the robustness of model evaluation in video-derived frames datasets.
II Proposed Method
An overview of the proposed framework is presented in Figure 1. The process begins by considering a collection of unlabeled videos denoted as $\mathcal{V}=\{V_{1},V_{2},\ldots,V_{K}\}$, where $K$ represents the total number of videos in the dataset. Each video $V_{k}$ is decomposed into a sequence of frames $\{I_{k,1},I_{k,2},\ldots,I_{k,N_{k}}\}$, where $N_{k}$ denotes the number of frames extracted from video $V_{k}$.
Figure 1: Illustration of the proposed cluster-based frame selection pipeline. Each video is decomposed into individual frames, from which features are extracted using semantic representations (e.g., CLIP [10]), handcrafted descriptors (e.g., HOG [3]), or lightweight learned features (e.g., XFeat [9]). Dimensionality reduction (e.g., PaCMAP [16]) and clustering (e.g., HDBSCAN [8]) are then applied to group visually similar frames before dataset partitioning.
II-A Feature Extraction
For each frame $I_{k,i}$, a feature vector $\mathbf{f}_{k,i}\in\mathbb{R}^{d}$ is extracted using handcrafted or learned descriptors. Classical descriptors such as the scale-invariant feature transform (SIFT) [7, 6] and the histogram of oriented gradients (HOG) [3] capture local texture and shape information by encoding gradient orientation patterns within localized image regions.
The Accelerated Features for Lightweight Image Matching (XFeat) [9] descriptor provides efficient keypoint detection and description through a lightweight convolutional architecture, while deep pretrained models such as Contrastive Language-Image Pretraining (CLIP) [10], Sigmoid Loss for Language Image Pre-training (SigLIP) [17], and DINO-V3 [13] offer semantic image representations that can capture visual concepts. The Vector of Locally Aggregated Descriptors (VLAD) [5] was applied to both SIFT and XFeat to combine their local keypoint descriptors into a single compact, fixed-length feature vector. VLAD creates a compact representation by assigning each local descriptor to its nearest codeword in a learned codebook and aggregating the residuals between them into a single vector. The feature extraction process is defined as:
$\mathbf{f}_{k,i}=\Phi_{\text{feat}}(I_{k,i}),$ (1)
where $\Phi_{\text{feat}}(\cdot)$ denotes the chosen feature extraction function.
II-B Dimensionality Reduction and Clustering
The extracted feature vectors $\mathbf{f}_{k,i}$ are projected into a low-dimensional embedding space using the pairwise controlled manifold approximation projection (PaCMAP) algorithm [16]:
$\mathbf{z}_{k,i}=\mathcal{P}_{\text{PaCMAP}}(\mathbf{f}_{k,i})$ (2)
where $\mathbf{z}_{k,i}\in\mathbb{R}^{m}$ represents the embedded frame in the latent space, and $\mathcal{P}_{\text{PaCMAP}}(\cdot)$ denotes the projection operator applied by the PaCMAP algorithm. Subsequently, the embedded representations $\mathbf{z}_{k,i}$} are clustered using the hierarchy of density-based spatial clustering (HDBSCAN) [8] algorithm. Due to the continuous and non-uniform nature of video data, density-based clustering methods such as HDBSCAN are generally more suitable than centroid-based approaches (e.g., K-Means [1] algorithm), as they can capture irregular cluster shapes and variable frame densities, well-suited for this task.
II-C Cluster-Based Dataset Partitioning
The HDBSCAN clustering algorithm is employed to group similar frames based on their feature representations $\mathbf{z}_{k,i}$. Let $C_{j}$ denote the $j$-th cluster containing feature $\mathbf{z}_{k,i}$, which correspond to frame $I_{k,i}$. Each cluster $C_{j}$ represents visually related frames, which are then assigned to the same dataset partition to prevent data leakage between sets.
III Experiments and Results
We analyze the results in Table I using the validation set of the ImageNet-VID set of the ImageNet Large Scale Visual Recognition Challenge 2015 (ILSVRC2015) [12] and all partitions of the UCF101 [14] dataset. The ImageNet dataset provides annotated images categorized by object synsets, whereas the UCF101 dataset consists of trimmed video clips labeled at the video level. To reduce visual redundancy and ensure that consecutive frames are not nearly identical, we extracted one frame per second from each video in the UCF101 dataset. These structured annotations allow us to detect and mitigate potential data leakage by ensuring that all frames from a given video remain within a single cluster. In our experiments, the images were resized to $224{\times}224$ for XFeat, CLIP, DINO, and SigLIP, and to either $128{\times}128$ or $224{\times}224$ for HOG, with the $128{\times}128$ setting performing slightly better while also producing a lower-dimensional descriptor. The VLAD vectors were reduced to $1024$ dimensions to provide a uniform feature representation, and the PaCMAP embeddings $\mathbf{z}_{k,i}$ were projected to a 256-dimensional space ($m=256$) to ensure a consistent low-dimensional representation across all feature extraction methods. To evaluate the effectiveness of the proposed frame grouping method, we assess clustering quality using the Adjusted Mutual Information (AMI) [15] and V-measure [11] metrics, which quantify how well the resulting clusters correspond to the ground-truth videos. The AMI measures the agreement between predicted clusters and true labels while accounting for chance, whereas the V-measure evaluates the trade-off between cluster homogeneity (the extent to which each cluster is composed of samples from a single class) and completeness (the extent to which samples of the same class share a cluster).
TABLE I: Clustering performance of HDBSCAN on validation sets from the ImageNet-VID and UCF101 datasets using different feature extractions and PaCMAP (256-D).
Feature Extraction Dataset V-measure AMI
SIFT + VLAD ImageNet-VID 0.81 0.80
UCF101 0.57 0.38
HOG ImageNet-VID 0.82 0.81
UCF101 0.61 0.48
HOG ($128\times 128$) ImageNet-VID 0.87 0.86
UCF101 0.67 0.54
XFeat + VLAD ImageNet-VID 0.90 0.89
UCF101 0.72 0.58
CLIP (ViT-B/32) ImageNet-VID 0.92 0.91
UCF101 0.75 0.66
SigLIP (ViT-B/16) ImageNet-VID 0.93 0.92
UCF101 0.75 0.67
DINO-V3 (ViT-B/16) ImageNet-VID 0.96 0.96
UCF101 0.87 0.80
In Table I, all feature extraction methods use an input image size of $224\times 224$ except for HOG ($128\times 128$). The results show that representations from deep pretrained models outperform classical and lightweight descriptors on both datasets, demonstrating their superior ability to capture the similarity of information leakage detection. In particular, DINO-V3 achieves the highest V-measure and AMI scores on both ImageNet-VID and UCF101. XFeat, combined with VLAD, also provides a noticeable improvement over traditional descriptors, while SIFT + VLAD remains the weakest across both datasets, especially on UCF101, where temporal variability increases the difficulty of clustering.
IV Conclusions and Future Work
This work presents a simple cluster-based frame selection strategy designed to minimize information leakage in video datasets. By grouping visually similar and temporally correlated frames before dataset partitioning, the proposed method can improve diversity and provide a fair evaluation setting, thereby mitigating overfitting in object detection models trained on video datasets. The experimental results demonstrate that DINO-V3 embeddings achieve the best clustering across both ImageNet-VID and UCF101 datasets, highlighting the advantage of representations for identifying information leakage. One limitation of the approach is its reliance on HDBSCAN and its hyperparameter choices, and we plan to explore adaptive clustering strategies that may address this limitation. We also aim to quantify the performance gap in image-based object detection models by training with and without the proposed approach to assess the impact of information leakage across different datasets.